The speedometer on a car is calibrated to read faster than the actual speed. A speed of 55km/h show as 60km/h on the speedometer

a) Calculate the actual speed when the speedometer is showing 90km/h?

b) What speed shows on the speedometer when the car is travelling at 120km/h?


1
Expert's answer
2022-04-08T02:37:48-0400

The calibration coefficient equals:


"k=\frac{60}{55}=1.09."


The actual speed when the speedometer is showing 90km/h equals:


"v_a=\frac{90}{k}=\frac{90}{1.09}=82.6\space km/h."


When the car is travelling at 120km/h, the speedometer shows:


"v_b=120k=120\cdot 1.09=130.8 \space km/h."